OpenStudy (grace111):

Help ----> medal . (1) - Centimetres are a measure of length . What are square centimeters a measure of ? (2) - Twenty - five percent of a number is 8 . What is the number ?

OpenStudy (anonymous):

1.Area

OpenStudy (anonymous):

2. i think its 32%

OpenStudy (anonymous):

1. Area 2. 40

OpenStudy (grace111):

thank you all :)

OpenStudy (anonymous):

welcome but 2nd one is 40 not 32 :)

OpenStudy (anonymous):

How did you get 40??

OpenStudy (anonymous):

We both get medals there xD

OpenStudy (anonymous):

XD

OpenStudy (grace111):

xD xD you two deserve them

OpenStudy (anonymous):

dats how to get 40 20 % = ( 8 / Y ) x 100 Y x 20 = 8 x 100 Y = 800 / 20 Y = 40

OpenStudy (anonymous):

and i dont help for medals